# An Empirical Study of Self-Admitted Technical Debt in Machine Learning Software

### v2

@@ 5.1. RQ1: What is the prevalence of SATD in ML-based systems?
Summary of RQ1 Machine learning projects have a median percentage of SATD that is twice the median percentage of SATD in non-machine learning projects. Overall, 3.5% of ML comments have SATD whereas only 2% of non-ML software has SATD. Such results are attributed to higher churn in ML projects.
@@ 5.2. RQ2: What are the different types of SATD in ML-based systems?
Figure 8. Distribution of SATD in the studied ML domains.
Table 6. Distribution of different types of SATD in ML software. Types except for the bolded ones are reused from Bavota and Russo (bavota2016large, ). The provided examples are hyperlinked to the source code file.
@@
On Hold - [D]: Existing SATD comments in the code even though the debt has been paid. [E]:“todo maxlen” (maximum length implementation was already done in code) 3.2
Unknown - [D]: Debt unclear from the description. [E]:“TODO{CJK*}UTF8gbsn 算法 4” 1.2
@@
In addition to the traditional SATD taxonomy by Bavota and Russo, ML code faces two new types of SATDs, namely configuration debt and inadequate tests debt. The most frequent type of SATD in ML code is functional requirement debt. Table 5.2 shows the distribution of SATDs observed in our studied sample. We found instances of code debt, design debt, requirement debt, and test debt, which are identified in the taxonomy proposed by Bavota and Russo (bavota2016large, ). Functional requirements debt accounts for 34% of SATD in our sample. Among them, 19% is related to features needing improvement (i.e., the SATD comment contains the statement “improvement of features needed”) and 15% of functional requirement debt is about new features (i.e., the SATD comment contain the statement “new features to be implemented”). Code SATD accounts for a total of 30% debt with low internal quality, refactoring and workaround SATDs, each accounting for 11%, 10% and 9% of the instances respectively. Known defect to fix accounts for 8% of SATDs instances, while SATDs related to design patterns account for 6.1% of SATDs instances. We also identified eight instances (or 3%) of On-Hold SATD. On-Hold SATDs are comments referring to a debt that has already been paid (Rungroj, ), but the SATD comment was not removed from the code. We identified two new categories of SATD not mentioned in the original taxonomy; i.e., configuration debt and inadequate tests, with 42 and 17 instances, respectively.
A visual representation of the debt taxonomy is provided in Figure 9, where the uncolored debt types were not found in our research. The newly found ML debt types are highlighted in the darker color palette in Figure 9, whereas the lighter orange colored ones are reused from Bavota et al.’s extended taxonomy (bavota2016large, ; Rungroj, ; suryanarayana2014refactoring, ).
@@
In our context, configuration debt refers to sub-optimal settings or lack of clear documentation on essential parameters governing ML system behavior, which includes, but is not limited to, algorithm hyperparameters, model settings, and data pipeline configurations. For instance, the comment “TODO Also add batch_size and other extraction specific configs to file [^16] Accessed on: 2024-03-22, https://github.com/Cloud-CV/visual-chatbot/tree/master//viscap/data/extract_features_maskrcnn.py signifies configuration debt as it alludes to an absence of standardized practices to track and manage essential configurations like batch sizes and extraction settings, potentially leading to inefficiencies and challenges in model reproducibility. Such configurations are needed to pre-process the data, as well as to tune the model (aka, hyper-parameters). Another example of such configuration debt is found in the project Image-Processing-via-deep-learning [^17] Accessed on: 2024-03-22, https://github.com/eddieczc/Image-Processing-via-deep-learning, where a developer mentioned that “300 iterations seems good enough but you can certainly train longer”. This SATD comment indicates the need to update hyperparameters’ configuration. Similarly, a developer [^18] Accessed on: 2024-03-22, https://mypublicdoc.s3.us-east-2.amazonaws.com/DebtFiles/24%23%5Eviscap%5Edata%5Eextract_features_maskrcnn.py mentioned “# TODO: Also add batch_size and other extraction specific configs to file”, indicating improper management of configuration variables. Regarding the inadequate test SATDs, they occur when testing is postponed to the future. For instance, in project [^19] Accessed on: 2024-03-22, https://github.com/NVIDIA/DeepLearningExamples.git, developer(s) mentioned “# TODO(yonib): Also test logit_scale with anchorwise=False” indicating missing / inadequate tests before committing the code. The taxonomy proposed by Bavota and Russo contains a “Test” category, but this category refers to the quality of test cases, while our new inadequate tests category is about missing or inadequate testing against a functional aspect of the code. The absence of a necessary test case extends “test debt” in its traditional scope, that presumes the existence of some testing but of insufficient quality. This distinction highlights the severe impact of missing tests on the functionality and reliability of software, particularly in ML systems where such omissions can be particularly detrimental.
@@
Bavota and Russo (bavota2016large, ) also found that instances of Code, Requirement, Defect, Design, Test, and Documentation debt account for respectively, 30%, 20%, 20%, 13%, 8%, and 7% of SATD in traditional software projects. In this study, we found them to represent respectively 30%, 36%, 8%, 13%, 5%, and 6% of SATDs in ML software projects (as shown in Figure 10). This indicates that ML software has lower proportions of defect and design debts and higher proportions of requirement debt than traditional (non-ML) software projects. However, the proportions of code, test, and documentation debts are similar, while we also found new SATD types of Configuration debt (12%) and Inadequate tests (8%).
@@
Figure 10. Comparison of types of debt identified by our research and prior research conducted by Bavota et al. (bavota2016large, ). On-Hold and Configuration debts are unique to ML-based systems.
@@ 5.3. RQ3: Which stages of the ML pipeline are more prone to SATD?
Model Building has the largest proportion of SATDs (46.7%), whereas Model Deployment has the lowest, i.e., 0.9%. This result is not very surprising given the amount of code churn that the Model Building stage experiences, as a result of the frequent retraining that is required to cope with concept drift issues.
@@ 5.4. RQ4: How long does SATD survive in ML-based systems?
Debt addition. SATDs are introduced in ML projects 140 times earlier during the development process in comparison with non-ML projects. A median ML debt is added in 10 days, while a median non-ML debt is added in 1,405 days. Figure 12 shows the survival curves in ML and non-ML code, with the difference between both curves is statistically significant with Wilcoxon rank sum test p-value=0.0. The steep decline of the ML curve in Figure 12 reflects a faster decline of the probability of ML projects to remain clean, i.e., remain without SATD. Within just one year of development time, ML projects have a low probability (i.e., 0.36) of its source code remaining debt-free, whereas non-ML projects have a much higher probability (i.e., 0.9). Even after 4 years, the probability of SATD introduction in non-ML projects is still below 0.3 (i.e., probability of SATD-free source code above 0.7), while it is 0.05 for ML projects.
Debt removal. SATD is removed 3.7 times faster in ML projects in comparison with non-ML projects. After its addition, a median ML SATD is removed in 1.1 years, whereas a non-ML SATD is removed in 4.1 years. Figure 13 shows the likelihood of debt removal in ML and non-ML projects. After a two-year time span, the probability of SATD to survive in ML projects is 0.89, while this probability is 0.94 for non-ML projects.
Such observations imply that ML developers may be cognizant of the faster debt accumulation in ML code and hence dedicate more time and effort towards its mitigation than developers working on non-ML code. Another reason could be that the shorter lifetime of SATD in ML projects may also be due to the shorter life cycle of these projects. In fact, in our data set, the medium lifecycle of ML projects is 1.15 years, while it is 4.16 years for non-ML projects.
Another possible reason could be the fact that the repetitive nature of ML experiments flushes out debt while dead experimental code paths are being refactored. This is corroborated by the 52% of cases of SATD removal in ML code that coincided with the deletion of the actual source code file. For the remaining 48% of cases, the file was only modified. This may indicate that SATD removal often happens during massive refactoring, where the debt-containing files are deleted.
@@
Overall, our findings corroborate the idea that SATD in ML projects is both introduced and resolved more rapidly than in non-ML projects. This higher code churn rate in ML software development was further discussed and evidenced in the results of RQ1 (see Section 5.1), providing additional support for our initial assertions. Prior research found that higher churn also correlates with higher likelihood of bugs (Morales15, ; mauerer2021search, ).
@@ 5.5. RQ5: What are the characteristics of long-lasting SATDs in ML-based systems?
Figure 15 shows mean Shapely values obtained for the best performing Random Forest model (from the bootstrapping analysis). Our models from the bootstrapping analysis have a median Precision of 93%, Recall of 98.8%, and F1 Score of 95.8%. SATD introduced during large code changes (Churn_[diff]) spanning multiple files (MD_[diff]) are likely to be long-lasting SATD according to our explanatory model. The timing of the commit, i.e., PT_[hist] (Project Time) and LCT_[hist] (Last Commit Time) also appear to be important factors in the duration of SATDs in the studied projects. Long-lasting SATDs are added at a median of 231 days, as compared to quickly removed SATDs which are added at a median of 7 days (i.e., towards the beginning of the project).
@@
Adding to the above hypothesis, it is also probable that resource allocation might be a factor, with developers focusing their efforts on managing SATD in larger, more critical files, similar to how testing larger files might find more bugs (mcintosh2014impact, ; zhang2012empirical, ). Risk perception could also play a part, as SATD in smaller files might be viewed as less likely to cause major issues. Familiarity might contribute too, as developers are potentially more at ease managing SATD in larger files they frequently work on. Oversight during code reviews, due to the less complex nature of small files, might allow more SATD to persist. Developers could also underestimate the impact of SATD in these files, falsely assuming its repercussions will be negligible. Additionally, it is plausible that smaller files, deemed easier to rewrite or refactor in the future, might have their SATD issues postponed. Lastly, the cognitive load to understand and fix SATD in larger files might draw developers’ focus away from smaller ones. Future work should validate these hypotheses.
@@ 6.1. Internal Validity
The act of “self-admitting” technical debt can manifest through various channels, including GitHub issues or project management tools such as Trello and Jira. Consequently, measuring SATD solely through code comments might not encompass the entire spectrum of ”self-admission” of debt in software projects. Nevertheless, Potdar and Shihab (potdar2014exploratory, ) have highlighted SATD in code comments as a discernible metric for (SA)TD. Furthermore, we follow authoritative research ((liu2018satd, ; Zampetti, ; liu2020using, ; alves2014towards, ; maldonado2015detecting, ; bavota2016large, ; da2017using, ; wattanakriengkrai2018identifying, ; liu2020using, ; kruchten2012technical, ; kruchten2013technical, )) in the area of self-admitted TD, that have estimated SATD via code comments.
The SATD identification tool used in our study may provide false positives. Although there are many SATD identification tools, we chose a NLP-based SATD detector proposed by Liu et al. (liu2018satd, ). This tool, unlike other state-of-the-art debt detector tools that provide high precision and low recall, has a high recall. Moreover, this tool has been utilized in several previous studies, such as those by Liu et al. (liu2021exploratory, ) and Zampetti et al. (Zampetti, ).
In addressing RQ2 and RQ3, two of the authors manually examined 611 instances of SATDs. This process may have introduced a subjective bias. However, given the high Cohen Kappa score of 94% between the two authors, we believe that the risk of bias is minimal. Furthermore, to compare the qualitative results of RQ2, we rely on the findings of Bavota and Russo (bavota2016large, ) and do not use our sample of 318 non-ML projects. This is because our extensive qualitative analysis, which took several months of efforts to label the 611 SATD samples, is non-trivial. Replicating such efforts on a sample of non-ML SATD instances was not possible in the scope of this study.

### v3

@@ 5.2. Discussion: Explanatory analysis for RQ1
Since ML developers tend to experiment frequently to find the optimal model, the ML codebase inherently undergoes significant modifications. Given that high code churn is widely recognized as a risk factor associated with poorer software quality and higher technical debt in general software engineering literature (mauerer2021search; Morales15), we hypothesize it could similarly influence SATD prevalence in ML contexts. To investigate this explanatory hypothesis formally, we state the hypotheses:
H0: There is no significant correlation between code churn, file changes, file complexity, and the proportion of SATD in ML and non-ML projects.
As mentioned in 3, we adopt Devil’s Advocate approach, individually testing each potential confound—churn, file changes, and complexity to see if any of them alone can explain away the ML vs non-ML SATD gap.
Results. Our analysis shows (Figure 7(a)) that ML code indeed experiences significantly higher code churn compared to non-ML code (Wilcoxon Rank Sum test $p<0.01$, Cliff’s $\delta=0.8$), aligning with our expectation from the repetitive and iterative nature of ML development (e.g., continuous hyperparameter tuning and dataset updates). Additionally, SATD-containing files in ML code exhibit significantly higher churn normalized by file size (Figure 7(b)), suggesting churn as a notable risk factor for SATD occurrence in ML contexts.
To test the devil’s advocate, i.e., to rule out alternative explanations, we also analyze two other potential confounding factors—file-level complexity and the number of files changed per commit (diffusion). Contrary to churn, we find that file complexity is significantly lower in ML compared to non-ML projects (Figure 6(b); Wilcoxon Rank Sum test $p<0.01$, Cliff’s $\delta=0.2$ for SATD files), and the number of files changed per commit also shows negligible differences (Figure 6(a); Cliff’s $\delta<0.1$). These results indicate that neither complexity nor the number of files changed per commit significantly explains the observed higher SATD prevalence in ML projects.
Overall, after explicitly controlling for potential confounding factors (code churn, file complexity, and file-change diffusion), the significantly higher prevalence of SATD in ML code persists. This strongly suggests that the high SATD density observed in ML projects is not merely an artifact of these structural metrics, but rather intrinsically linked to the specific nature and iterative development processes common in ML software development. ML developers should be aware of this intrinsic risk and proactively manage their technical debt, potentially through systematic monitoring of SATD indicators such as churn during model building or frequent hyperparameter updates.
Figure 8. Distribution of SATD in the studied ML domains.
@@ 5.3. RQ2: What are the different types of SATD in ML-based systems?
Table 6. Distribution of different types of SATD in ML software. Types except for the bolded ones are reused from Bavota and Russo (bavota2016large). The provided examples are hyperlinked to the source code file.
@@
On-Hold - [D]: Existing SATD comments in the code even though the debt has been paid. [E]:“todo maxlen” (maximum length implementation was already done in code) 3.2
Unknown - [D]: Debt unclear from the description. [E]:“TODO 算法 4” 1.2
@@
In addition to the traditional SATD taxonomy by Bavota and Russo, ML code faces two new types of SATDs, namely configuration debt and inadequate tests debt. The most frequent type of SATD in ML code is functional requirement debt. Table 6 shows the distribution of SATDs observed in our studied sample. We found instances of code debt, design debt, requirement debt, and test debt, which are identified in the taxonomy proposed by Bavota and Russo (bavota2016large). Functional requirements debt accounts for 34% of SATD in our sample. Among them, 19% is related to features needing improvement (i.e., the SATD comment contains the statement “improvement of features needed”) and 15% of functional requirement debt is about new features (i.e., the SATD comment contain the statement “new features to be implemented”). Code SATD accounts for a total of 30% debt with low internal quality, refactoring and workaround SATDs, each accounting for 11%, 10% and 9% of the instances respectively. Known defect to fix accounts for 8% of SATDs instances, while SATDs related to design patterns account for 6.1% of SATDs instances. We also identified eight instances (or 3%) of On-Hold SATD. On-Hold SATDs are comments in which developers explicitly document that implementation is paused pending an external condition or event (e.g., an open issue must be resolved or a dependency implemented elsewhere) (maipradit2020wait; Rungroj), yet the SATD comment remains in the code until that condition is met. We identified two new categories of SATD not mentioned in the original taxonomy; i.e., configuration debt and inadequate tests, with 42 and 17 instances, respectively.
A visual representation of the debt taxonomy is provided in Figure 9, where the uncolored debt types were not found in our research. The newly found ML debt types are highlighted in the darker color palette in Figure 9, whereas the lighter orange colored ones are reused from Bavota et al.’s extended taxonomy (bavota2016large; Rungroj; suryanarayana2014refactoring).
@@
In our context, configuration debt refers to sub-optimal settings or lack of clear documentation on essential parameters governing ML system behavior, which includes, but is not limited to, algorithm hyperparameters, model settings, and data pipeline configurations. For instance, the comment “TODO Also add batch_size and other extraction specific configs to file [^18] Accessed on: 2024-03-22, https://github.com/Cloud-CV/visual-chatbot/tree/master//viscap/data/extract_features_maskrcnn.py signifies configuration debt as it alludes to an absence of standardized practices to track and manage essential configurations like batch sizes and extraction settings, potentially leading to inefficiencies and challenges in model reproducibility. Such configurations are needed to pre-process the data, as well as to tune the model (aka, hyper-parameters). Another example of such configuration debt is found in the project Image-Processing-via-deep-learning [^19] Accessed on: 2024-03-22, https://github.com/eddieczc/Image-Processing-via-deep-learning, where a developer mentioned that “300 iterations seems good enough but you can certainly train longer”. This SATD comment indicates the need to update hyperparameters’ configuration. Similarly, a developer [^20] Accessed on: 2024-03-22, https://mypublicdoc.s3.us-east-2.amazonaws.com/DebtFiles/24%23%5Eviscap%5Edata%5Eextract_features_maskrcnn.py mentioned “# TODO: Also add batch_size and other extraction specific configs to file”, indicating improper management of configuration variables.
Regarding the inadequate test SATDs, they occur when testing is postponed to the future. For instance, in project [^21] Accessed on: 2024-03-22, https://github.com/NVIDIA/DeepLearningExamples.git, developer(s) mentioned “# TODO(yonib): Also test logit_scale with anchorwise=False” indicating missing / inadequate tests before committing the code. The taxonomy proposed by Bavota and Russo contains a “Test” category, but this category refers to the quality of test cases, while our new inadequate tests category is about missing or inadequate testing against a functional aspect of the code. The absence of a necessary test case extends “test debt” in its traditional scope, that presumes the existence of some testing but of insufficient quality. This distinction highlights the severe impact of missing tests on the functionality and reliability of software, particularly in ML systems where such omissions can be particularly detrimental.
@@
Bavota and Russo (bavota2016large) also found that instances of Code, Requirement, Defect, Design, Test, and Documentation debt account for respectively, 30%, 20%, 20%, 13%, 8%, and 7% of SATD in traditional software projects. In this study, we found them to represent respectively 30%, 36%, 8%, 13%, 5%, and 6% of SATDs in ML software projects (as shown in Figure 10). This indicates that ML software has lower proportions of defect and design debts and higher proportions of requirement debt than traditional (non-ML) software projects. However, the proportions of code, test, and documentation debts are similar, while we also found new SATD types of Configuration debt (12%) and Inadequate tests (8%).
@@
Figure 10. Comparison of types of debt identified by our research and prior research conducted by Bavota et al. (bavota2016large). On-Hold and Configuration debts are unique to ML-based systems.
@@ 5.4. RQ3: Which stages of the ML pipeline are more prone to SATD?
Model Building has the largest proportion of SATDs (46.7%), whereas Model Validation has the lowest, i.e., 0.4%. This result is not very surprising given the amount of code churn that the Model Building stage experiences, as a result of the frequent retraining that is required to cope with concept drift issues.
@@ 5.5. RQ4: How long does SATD survive in ML-based systems?
Debt addition. SATDs are introduced in 140 times earlier during the development process of the studied ML projects in comparison with non-ML projects. A median ML debt is added in 10 days, while a median non-ML debt is added in 1,405 days. Figure 12 shows the survival curves in ML and non-ML code, with the difference between both curves is statistically significant with Wilcoxon rank sum test p-value=0.0. The steep decline of the ML curve in Figure 12 reflects a faster decline of the probability of ML projects to remain clean, i.e., remain without SATD. Within just one year of development time, ML projects have a low probability (i.e., 0.36) of its source code remaining debt-free, whereas non-ML projects have a much higher probability (i.e., 0.9). Even after 4 years, the probability of SATD introduction in non-ML projects is still below 0.3 (i.e., probability of SATD-free source code above 0.7), while it is 0.05 for ML projects.
Debt removal. The survival analysis indicates 3.7 times faster SATD removal in ML projects in comparison with non-ML projects in our studied dataset. After its addition, a median ML SATD is removed in 1.1 years, whereas a non-ML SATD is removed in 4.1 years. Figure 13 shows the likelihood of debt removal in ML and non-ML projects. After a two-year time span, the probability of SATD to survive in ML projects is 0.89, while this probability is 0.94 for non-ML projects. Such observations imply that ML developers may be cognizant of the faster debt accumulation in ML code and hence dedicate more time and effort towards its mitigation than developers working on non-ML code.
While our findings corroborate the idea that SATD in ML projects is both introduced and resolved more rapidly than in non-ML projects, it is important to put these numbers in context. Alternative reasons include 1) shorter lifetime of SATD in ML projects may be due to the shorter life cycle of these projects. In fact, in our data set, the medium lifecycle of ML projects is 1.15 years, while it is 4.16 years for non-ML projects. This is expected, as ML software was relatively new when the data for this study was collected; 2) the repetitive nature of ML experiments may flush out debt while dead experimental code paths are being refactored. This is corroborated by the 52% of cases of SATD removal in ML code that coincided with the deletion of the actual source code file. For the remaining 48% of cases, the file was only modified. This may indicate that SATD removal often happens during massive refactoring, where the debt-containing files are deleted. 3) the high code churn characteristic of ML projects, reflecting their fast-paced and iterative development cycles. This higher code churn rate in ML software development was further discussed and evidenced in the results of RQ1 (see Section 5.1) and Figure (4-d), providing additional support for our initial assertions. Furthermore, prior research found that higher churn also correlates with higher likelihood of bugs (Morales15; mauerer2021search).
Overall, further research, particularly qualitative studies are needed to determine whether rapid SATD turnover in ML projects reflects healthy refactoring or risk-prone instability.
@@ 5.6. RQ5: What are the characteristics of long-lasting SATDs in ML-based systems?
Figure 15 shows mean Shapely values obtained for the best performing Random Forest model (from the bootstrapping analysis). Our models from the bootstrapping analysis have a median Precision of 93%, Recall of 98.8%, and F1 Score of 95.8%.
SATD introduced during large code changes (Churn_[diff]) spanning multiple files (MD_[diff]) are likely to be long-lasting SATD according to our explanatory model. The timing of the commit, i.e., PT_[hist] (Project Time) and LCT_[hist] (Last Commit Time) also appear to be important factors in the duration of SATDs in the studied projects. Long-lasting SATDs are added at a median of 231 days, as compared to quickly removed SATDs which are added at a median of 7 days (i.e., towards the beginning of the project).
@@
Adding to the above hypothesis, it is also probable that resource allocation might be a factor, with developers focusing their efforts on managing SATD in larger, more critical files, similar to how testing larger files might find more bugs (mcintosh2014impact; zhang2012empirical). Risk perception could also play a part, as SATD in smaller files might be viewed as less likely to cause major issues. Familiarity might contribute too, as developers are potentially more at ease managing SATD in larger files they frequently work on. Oversight during code reviews, due to the less complex nature of small files, might allow more SATD to persist. Developers could also underestimate the impact of SATD in these files, falsely assuming its repercussions will be negligible. Additionally, it is plausible that smaller files, deemed easier to rewrite or refactor in the future, might have their SATD issues postponed. Lastly, the cognitive load to understand and fix SATD in larger files might draw developers’ focus away from smaller ones. Future work should validate these hypotheses.
@@ 6.1. Internal Validity
The act of “self-admitting” technical debt can manifest through various channels, including GitHub issues or project management tools such as Trello and Jira. Consequently, measuring SATD solely through code comments might not encompass the entire spectrum of ”self-admission” of debt in software projects. Nevertheless, Potdar and Shihab (potdar2014exploratory) have highlighted SATD in code comments as a discernible metric for (SA)TD. Furthermore, we follow authoritative research ((liu2018satd; Zampetti; liu2020using; alves2014towards; maldonado2015detecting; bavota2016large; da2017using; wattanakriengkrai2018identifying; liu2020using; kruchten2012technical; kruchten2013technical)) in the area of self-admitted TD, that have estimated SATD via code comments.
The SATD identification tool used in our study may provide false positives. Although there are many SATD identification tools, we chose a NLP-based SATD detector proposed by Liu et al. (liu2018satd). This tool, unlike other state-of-the-art debt detector tools that provide high precision and low recall, has a high recall. Moreover, this tool has been utilized in several previous studies, such as those by Liu et al. (liu2021exploratory) and Zampetti et al. (Zampetti).
In addressing RQ2 and RQ3, two of the authors manually examined 611 instances of SATDs. This process may have introduced a subjective bias. However, given the high Cohen Kappa score of 94% between the two authors, we believe that the risk of bias is minimal. Furthermore, to compare the qualitative results of RQ2, we rely on the findings of Bavota and Russo (bavota2016large) and do not use our sample of 318 non-ML projects. This is because our extensive qualitative analysis, which took several months of efforts to label the 611 SATD samples, is non-trivial. Replicating such efforts on a sample of non-ML SATD instances was not possible in the scope of this study.